The future value of an investment is the value of the investment after a certain period of time and is given by the formula: FV=PV(l+i)^n in which
FV= future value, PV= present value, i=the interest rate, n= the number of years invested
Find the future value of an investment of $120,000 in 5 years from now if an investment earns 5% annual interest for the next 10 years.

Answer by nerdybill(2695)

Simply plug in what was given and solve:
FV=PV(l+i)^n
The problem gave us the following:
PV = 120000
i = .05
n = 10
.
FV=120000(l+.05)^10
FV=120000(l.05)^10
FV=120000(1.629)
FV=$195,467.36